OpenStudy (anonymous):

What is the value of 2|6 + p| - 3p3 when p = -2?

OpenStudy (anonymous):

This would be my guess (2|6 +(-2)|) - (3*-2*3) = 2| 3 | - (-27) = 6 + 27 = 33

OpenStudy (anonymous):

thanks man